the coefficient of x^n in the expa. of
(1+x+x^2+....)^(-n) is

Grade:12

1 Answers

Sona Srivastava
13 years ago

1 + x + x..... = 1/(1-x) .......infinite gp

(1/(1-x))-n = (1-x)n

solve ahead accordingly